A card is chosen at random from a deck of 52 cards. It is then replaced and a second card is chosen. What is the probability of choosing a jack and then an eight?


1
Expert's answer
2021-06-16T03:51:08-0400

Р(getting jack on first chance) "= \frac{4}{52}"

P(getting an eight on second chance) "= \frac{4}{52}"

P(choosing a jack and then an eight) "=\frac{4}{52} \times \frac{4}{52}"

"= \frac{1}{169}"